# From Planetesimals to Dwarf Planets by Pebble Accretion

### v1

@@ 2.2 Growth of bodies by pebble accretion
Figure 4 presents the Safronov and settling growth rates as functions of mass. Here, the rates are expressed in terms of the growth timescale, $t_{\mathrm{gr}}=m/(dm/dt)$, which is given in units of $t_{0}$ (see Equation 5). At low masses, accretion is dominated by gravitational focusing (Safronov limit; yellow curve). It is well-known that in this limit, the growth timescale decreases with the mass: the runaway growth regime (Wetherill & Stewart, 1989). Pebble accretion (the settling mechanism) is unimportant at low masses. Yet at a mass scale that is a fraction of $q_{\ast}=\sigma^{3}\tau_{s}$ (the mass corresponding to $\sigma\sim v_{\ast}$), the settling mechanism starts to become efficient, increasing the growth rate significantly (Visser & Ormel, 2016). In the case of a turbulence-dominant velocity field, this transition is slightly less abrupt (thin line). Hence, bodies more massive than ${\sim}0.1\,m_{\ast}$ will undergo rapid growth, while those with $m\ll m_{\ast}$ grow more slowly. Pebble accretion therefore has the potential to flatten the high-mass tail of the size distribution, illustrated in Figure 4 by the gray arrows.
@@ 2.3 Model parameters and likelihood
An overview of the parameters in the model is given in Table 1. The initial distribution of bodies is specified by parameters $m_{\mathrm{brk}}$ and the initial mass of the belt $m_{\mathrm{imf}}$ (see Equation 2), which are uniformly sampled in the log space. Pebble properties are given by their dimensionless velocity $\sigma$, the dimensionless stopping time $\tau_{s}$, and the total accreted pebbles mass $m_{\mathrm{Peb}}$. In addition to the continuous sampling of these parameters, the model also employs a binary switch for the nature of the velocity field laminar or turbulent. If the velocity is laminar, $\sigma$ is identified with the normalized pressure gradient parameter $\eta$; if turbulent, with the Mach number of the turbulent velocity field. The velocity field matters in calculating the $f_{\mathrm{set}}$ modulation factors (see Equation 11).
In total, the sampling provides us with several hundreds bodies, which covers the entire mass distribution. The mass of each of the bodies is integrated over time. We work in dimensionless units where time is normalized by $t_{0}$ (Equation 5) and mass by the mass of the Sun. The pebble accretion rate $dq/dt$ is computed for each of these bodies, according to the collision rates given in Section 2.2. The bodies do not interact among themselves; and because $dq/dt$ is an increasing function of $q$, the initial order of their masses stays preserved. Under these conditions, the calculation boils down to a set of ordinary differential equations, which we solve with the scipy.integrate.odeint library. Simultaneously with the increase of $q$, we decrease the mass of the pebble reservoir $q_{\mathrm{peb}}$, whose initial value is also a parameter. When $q_{\mathrm{peb}}$ reaches 0, we stop the simulation and record the elapsed time $t$ (in units of $t_{0}$). This distribution is then weighted by the implantation factor $f_{\mathrm{impl}}$, the ratio between numbers of bodies in the present-day P48 and in the primordial belt, to obtain the simulated distribution of bodies in P48.
@@
The likelihood function is calculated by comparing the simulated distribution with the reconstructed distribution in the OSSOS++ domain and the observed number of bodies in the MPC domain (see Figure 2). To this effect, the particles are binned by mass, with 10 bins covering the OSSOS++ region and 15 covering the MPC region (the bin width is about 0.2 dex). The number of bodies in the synthetic and observed distributions are then compared for each bin. In the OSSOS++ region, Gaussian statistics are assumed, i.e., the contribution to the log-likelihood of bin $k$ equals:
@@
$\ln P_{k}^{\mathrm{MPC}}=\ln\mathcal{B}(N_{k},N_{\mathrm{sim},k},f_{\mathrm{impl}})$ (13)
where $\mathcal{B}(x,n,p)$ is the probability of drawing $x$ successes out of $n$ independent samples where each draw has probability $p$.
In addition, we apply two priors to the likelihood function, related to the following observations:
  • The number of Jupiter Trojans constrain the total mass of the primordial belt $m_{\mathrm{imf}}+m_{\mathrm{peb}}$. Since the Jovian Trojans are also thought to be captured from the same disk, the mass of bodies present in the primordial belt can be estimated using a numerically simulated implantation rate and the observed mass of the Trojans (see Figure 3 and Nesvorný et al. 2013). The total pre-instability mass of the primordial belt falls in the range of $14{-}28\,m_{\oplus}$. Consequently, we add a term $\mathcal{P}^{\mathrm{PB}}$ to the likelihood function corresponding to a situation where the parameter $\mu=\ln(m_{\mathrm{imf}}+m_{\mathrm{Peb}})$ is normally distributed, centered around $\ln 20\,m_{\oplus}$ and with spread $\sigma=0.34$.
  • The preservation of the dynamically cold classicals, limits the mass of individual bodies. If any planetary-mass bodies formed in the primordial outer disk and were scattered out by a migrating Neptune, the cold classical belt will likely be overheated or even destroyed by the gravitational sweeping of such a planet. Therefore, the masses of the individual bodies making up the primordial belt could not have been too large. Petit et al. (1999) simulated the effect of massive protoplanets on the stability of the cold belt, and found a threshold mass of ${\sim}1\,m_{\oplus}$. To incorporate this mass constraint, we add a term $\ln\mathcal{P}^{\mathrm{CC}}=-(m_{1}/m_{\oplus})^{2}$ to the log-likelihood function, which suppresses models favoring formation of bodies beyond 1 Earth mass.
The total log-likelihood is then
$\ln\mathcal{P}=\sum_{k}\ln\mathcal{P}_{k}^{\mathrm{MPC}}+\sum_{k}\ln\mathcal{P}_{k}^{\mathrm{MPC}}+\ln\mathcal{P}^{\mathrm{CC}}+\ln\mathcal{P}^{\mathrm{PB}}$ (14)
We then conduct an MCMC analysis using emcee (Foreman-Mackey et al., 2013), marginalizing over the input parameters listed in Table 1.
Figure 5: MCMC posterior distribution of the single particle laminar model (lami-1p). The 16%, 50%, and 84% confidence levels are indicated above each histogram. Runs cluster around the line of constant $\tau_{s}\sigma^{3}$, which indicates a mass scale above which pebble accretion operates (panel b). High-velocity solutions, like in a smooth disk ($\sigma=\eta_{\mathrm{irr}}$), have difficulty fitting the data (panel a). The amount of accreted pebbles is lower than the mass of the initial popluation by around 0.4 dex (panel n).
@@ 3.1 Simulations with a single particle component
In physical units, $m_{\ast}=10^{C_{\tau\sigma}}m_{\odot}\approx 1.6\times 10^{23}\,\mathrm{g}$ (or $10^{-5}\,m_{\oplus}$; the actual mass where pebble accretion starts to dominate over Safronov accretion is ${\approx}10\%$ of this number; see Figure 4). For bodies of mass $m>m_{\ast}$, pebble accretion is fully operational, whereas below $m_{\ast}$ pebble accretion is suppressed. Therefore, for $m\ll m_{\ast}$, the mass distribution tends to follow the IMF, as given by the $m_{\mathrm{brk}}$ and $m_{\mathrm{imf}}$ parameters. However, due to the exponential tapering at high mass, the IMF cannot fit the high-mass end of the P48 bodies. Here, pebble accretion acts to flatten the mass distribution (see Figure 4). Table 2 shows that about $5{-}8$ Earth mass in pebbles are needed—an amount that is similar to the mass in the IMF.
@@
Figure 7: Final vs initial mass for several simulation runs (see Table 3). Vertical lines denote 10% of $m_{\ast}=\sigma^{3}\tau_{s}m_{\odot}$, which is the scale where pebble accretion kicks in. Dashed line is the identity line ($y=x$). Bodies ${\gtrsim}10^{22}$ grow their mass by pebble accretion many tenfolds and become dwarf planets. The gain in mass at low mass in the lami-1p-etairr and turb-1p-best is due to the slow pace of pebble accretion, allowing Safronov interactions to become important at low planetesimal mass.
@@
At high velocity, the correlation starts to break down. Here, the combination of high $\sigma$ and small $\tau_{s}$ renders pebble accretion very slow (Equation (10)) and allows ballistic encounters (Safronov focusing) to rival pebble accretion. As a result, the mass distribution at low mass is altered, which is demonstrated by the lami-1p-etairr model, which features a $\sigma$ value consistent with that of an irradiated disk. Figure 7 illustrates that in this model low-mass bodies accrete a significant amount of pebbles. As $\sigma$ increases, this trend becomes more pronounced resulting in poorer fits, because the OSSOS++ part of the distribution can no longer be matched. This finding is unsurprising as the shape of the IMF was adopted from the cold classicals, which was already shown to fit the shape of hot classicals across the OSSOS++ range (Kavelaars et al., 2021; Petit et al., 2023). But it also shows that accretion processes cannot maintain the shape of the low-mass part—at least not in conjunction with also matching the high-mass tail. It further lends support to the primordial nature of the low-mass part of the TNO size distribution.
@@
An additional concern for these small $\tau_{s}$ runs is that they take a long time ($t/t_{0}$ is high). We can estimate the characteristic accretion time for pebble accretion from the pebble accretion rate (Equation (10)): $t\sim q/\dot{q}\sim t_{0}/\tau_{s}$, for $f_{\mathrm{set}}\sim 1$ (i.e., $m>m_{\ast}$). This estimate agrees with Table 3 where we compiled $t/t_{0}$ for selected model runs. Recall that the simulations are conducted in dimensionless parameter space, where we can hide our ignorance of, e.g., the pebble density, which sets the value of $t_{0}$ (Equation 5). Yet, the pebble accretion mechanism requires gas-rich conditions and it must terminate before the lifetime of the gas-rich disk, typically estimated at ${\sim}5\,\mathrm{Myr}$ (Weiss et al., 2021). Therefore, a high $t/t_{0}$ renders these models implausible. We return to the timescale constraints in Section 4.
@@
Table 2 shows that a turbulent velocity field marginally improves the log-likelihood of the best model by an amount of +0.5. This specific run is characterized by a very low $\tau_{s}$ value (turb-1p-best entry in Table 3), which is disfavored in the light of above discussion. A higher $\tau_{s}$ run, following the $\tau_{s}$$\sigma$ correlation (turb-1p-hitau) fits the data equally well. Figure 6a presents cumulative distribution of several individual runs. All models overshoot the infliction point at $m=10^{-4}\,m_{\oplus}$ and also tend to undershoot the final bin where the observed CMF terminates (with Triton) at $m=3\times 10^{-3}\,m_{\oplus}$. These apparent mismatches are a consequence to the low number of TNOs at these masses, which can be understood from inspecting the differential distribution (on which the likelihood calculations are based). Figure 6b presents these fits for the MPC part of the distribution. Here, the error bars denote Poisson counting uncertainties. The bin to the right of $m=10^{-4}\,m_{\oplus}$ with 0 bodies in the MPC is typically fitted with ${\approx}2$ bodies in the simulation, but this still has a Poisson likelihood of 13.5%. Similarly, the simulation runs are quite insensitive to the presence of the two most massive bodies (Pluto and Triton); statistically, the low number of high-$m$ bodies do not carry much weight. As a result the runs that undershoot the cumulative curve in Figure 6a (e.g., turb-1p-best) statistically cannot be distinguished from those runs that include a larger $\tau_{s}$ (e.g., turb-1p-hitau) and those runs that contain more pebbles (e.g., turp-1p-himass).
@@ 3.2 A second particle component
Addition of a secondary component does not significantly improve the results. From Table 2, the Bayes factor (the difference in $\ln\mathcal{P}$) is $0.54$ between the laminar models and $0.18$ between the turbulence models. These values indicate that the added parameters have insignificant effects on the fits (Jeffreys, 1998). This insensitivity is reflected in Table 2 by the proximity of the size of the secondary component ($\tau_{2}$) to that of the main component ($\tau_{s}$) in the lami-2p and turb-2p models. Figure 6 presents this result graphically. The lami-2p-best run—the run where $\mathcal{P}$ is the highest—hardly stands out from the other runs. Curiously, these models tend to be fitted by low $\tau_{s}$ (and $\tau_{2}$) and correspondingly high $t_{0}$, indicating these runs improve the fit for the OSSOS++ part of the distribution, rather than the high-mass MPC part. In other words, the low number of high-mass bodies in P48 prevents any meaningful fit for the high-mass end of the distribution, which makes the model insensitive to a second particle component.
While we do not find evidence for a binary distribution of pebbles, two points should be noted. First, the size distribution could still be a power-law, which would ameliorate the $f_{\mathrm{set}}$ transition in a similar way as the turbulence velocity field would do (Lyra et al., 2023). Second, the lack of evidence for a second component does not preclude its presence. A bigger sample of high-mass bodies is necessary to establish or discount its existence.

### v2

@@ 2.2 Growth of bodies by pebble accretion
Figure 4 presents the Safronov and settling growth rates as functions of mass. Here, the rates are expressed in terms of the growth timescale, $t_{\mathrm{gr}}=m/(dm/dt)$, which is given in units of $t_{0}$ (see Equation 5). At low masses, accretion is dominated by gravitational focusing (Safronov limit; yellow curve). It is well known that in this limit, the growth timescale decreases with the mass: the runaway growth regime (Wetherill & Stewart, 1989). Pebble accretion (the settling mechanism) is unimportant at low masses. However, at a mass scale that is a fraction of $q_{\ast}=\sigma^{3}\tau_{s}$ (the mass corresponding to $\sigma\sim v_{\ast}$), the settling mechanism starts to become efficient, increasing the growth rate significantly (Visser & Ormel, 2016). In the case of a turbulence-dominant velocity field, this transition is slightly less abrupt (thin line). Hence, bodies more massive than ${\sim}0.1\,m_{\ast}$ will undergo rapid growth, while those with $m\ll m_{\ast}$ grow more slowly. Pebble accretion therefore has the potential to flatten the high-mass tail of the size distribution, illustrated in Figure 4 by the gray arrows.
@@ 2.3 Model parameters and likelihood
An overview of the parameters used in the model is given in Table 1. The initial distribution of bodies is specified by the parameter $m_{\mathrm{brk}}$ and the initial mass of the belt $m_{\mathrm{imf}}$ (see Equation 2), which are uniformly sampled in the log space. The pebble properties are given by their dimensionless velocity $\sigma$, the dimensionless stopping time $\tau_{s}$, and the total accreted pebbles mass $m_{\mathrm{Peb}}$. In addition to the continuous sampling of these parameters, the model also employs a binary switch for the nature of the velocity field, either laminar or turbulent. If the velocity is laminar, $\sigma$ is identified with the normalized pressure gradient parameter $\eta$; if the velocity is turbulent, $\sigma$ is identified with the Mach number of the turbulent velocity field. The velocity field matters in calculating the $f_{\mathrm{set}}$ modulation factors (see Equation 11).
In total, the sampling provides us with several hundred bodies, which covers the entire mass distribution. The mass of each of the bodies is integrated over time. We work in dimensionless units where time is normalized by $t_{0}$ (Equation 5) and mass by the mass of the Sun. The pebble accretion rate $dq/dt$ is computed for each of these bodies, according to the collision rates given in Section 2.2. The bodies do not interact among themselves, and because $dq/dt$ is an increasing function of $q$, the initial order of their masses stays preserved. Under these conditions, the calculation boils down to a set of ordinary differential equations, which we solve with the scipy.integrate.odeint library. Simultaneously with the increase in $q$, we decrease the mass of the pebble reservoir $q_{\mathrm{peb}}$, whose initial value is also a parameter. When $q_{\mathrm{peb}}$ reaches 0, we stop the simulation and record the elapsed time $t$ (in units of $t_{0}$). This distribution is then weighted by the implantation factor $f_{\mathrm{impl}}$, the ratio of the number of bodies in the present-day P48 to that in the primordial belt, to obtain the simulated distribution of bodies in P48.
@@
The likelihood function is calculated by comparing the simulated distribution with the reconstructed distribution in the OSSOS++ domain and the observed number of bodies in the MPC domain (see Figure 2). To this end, the particles are binned by mass, with 10 bins covering the OSSOS++ region and 15 covering the MPC region; the bin width is about 0.2 dex. The number of bodies in the synthetic and observed distributions are then compared for each bin. In the OSSOS++ region, Gaussian statistics are assumed; in other words, the contribution to the log-likelihood of bin $k$ equals
@@
$\ln P_{k}^{\mathrm{MPC}}=\ln\mathcal{B}(N_{k},N_{\mathrm{sim},k},f_{\mathrm{impl}}).$ (13)
Here $\mathcal{B}(x,n,p)$ is the probability of drawing $x$ successes out of $n$ independent samples where each draw has probability $p$.
In addition, we apply two priors to the likelihood function, related to the following observations. First, the number of Jupiter trojans constrains the total mass of the primordial belt $m_{\mathrm{imf}}+m_{\mathrm{peb}}$. Since the Jovian trojans are also thought to be captured from the same disk, the mass of bodies present in the primordial belt can be estimated using a numerically simulated implantation rate and the observed mass of the trojans (see Figure 3 and Nesvorný et al. 2013). The total pre-instability mass of the primordial belt falls in the range of $14{-}28\,m_{\oplus}$. Consequently, we add a term $\mathcal{P}^{\mathrm{PB}}$ to the likelihood function corresponding to a situation where the parameter $\mu=\ln(m_{\mathrm{imf}}+m_{\mathrm{Peb}})$ is normally distributed, centered around $\ln 20\,m_{\oplus}$ and with spread $\sigma=0.34$. Second, the preservation of the dynamically cold classicals limits the mass of individual bodies. If any planetary-mass bodies formed in the primordial outer disk and were scattered out by a migrating Neptune, the cold classical belt would likely be overheated or even destroyed by the gravitational sweeping of such a planet. Therefore, the masses of the individual bodies making up the primordial belt could not have been too large. Petit et al. (1999) simulated the effect of massive protoplanets on the stability of the cold belt, and found a threshold mass of ${\sim}1\,m_{\oplus}$. To incorporate this mass constraint, we add a term $\ln\mathcal{P}^{\mathrm{CC}}=-(m_{1}/m_{\oplus})^{2}$ to the log-likelihood function, which suppresses models favoring the formation of bodies beyond 1 Earth mass. The total log-likelihood is then
$\ln\mathcal{P}=\sum_{k}\ln\mathcal{P}_{k}^{\mathrm{MPC}}+\sum_{k}\ln\mathcal{P}_{k}^{\mathrm{MPC}}+\ln\mathcal{P}^{\mathrm{CC}}+\ln\mathcal{P}^{\mathrm{PB}}.$ (14)
We conduct an MCMC analysis using emcee (Foreman-Mackey et al., 2013), marginalizing over the input parameters listed in Table 1.
Figure 5: MCMC posterior distribution of the single particle laminar model (lami-1p). The 16%, 50%, and 84% confidence levels are indicated above each histogram. The runs cluster around the line of constant $\tau_{s}\sigma^{3}$, which indicates a mass scale above which pebble accretion operates (panel b). High-velocity solutions, as in a smooth disk ($\sigma=\eta_{\mathrm{irr}}$), have difficulty fitting the data (panel a). The mass of the accreted pebbles is lower than the mass of the initial popluation by around 0.4 dex (panel n).
@@ 3.1 Simulations with a single particle component
In physical units, $m_{\ast}=10^{C_{\tau\sigma}}m_{\odot}\approx 1.6\times 10^{23}\,\mathrm{g}$ (or $10^{-5}\,m_{\oplus}$; the actual mass where pebble accretion starts to dominate over Safronov accretion is ${\approx}10\%$ of this number; see Figure 4). For bodies of mass $m>m_{\ast}$, pebble accretion is fully operational, whereas below $m_{\ast}$ pebble accretion is suppressed. Therefore, for $m\ll m_{\ast}$, the mass distribution tends to follow the IMF, as given by the $m_{\mathrm{brk}}$ and $m_{\mathrm{imf}}$ parameters. However, due to the exponential tapering at high mass, the IMF cannot fit the high-mass end of the P48 bodies. Here pebble accretion acts to flatten the mass distribution (see Figure 4). Table 2 shows that about $5{-}8$ Earth masses in pebbles are needed, an amount that is similar to the mass in the IMF.
@@
Figure 7: Final vs initial mass for several simulation runs (see Table 3). The vertical lines denote 10% of $m_{\ast}=\sigma^{3}\tau_{s}m_{\odot}$, which is the scale where pebble accretion begins. The dashed line is the identity line ($y=x$). Bodies ${\gtrsim}10^{22}$ grow their mass by several tenfold by pebble accretion and become dwarf planets. The gain in mass at low mass in the lami-1p-etairr and turb-1p-best is due to the slow pace of pebble accretion, allowing Safronov interactions to become important at low planetesimal masses.
@@
At high velocity the correlation starts to break down. Here, the combination of high $\sigma$ and small $\tau_{s}$ renders pebble accretion very slow (Equation (10)) and allows ballistic encounters (Safronov focusing) to rival pebble accretion. As a result, the mass distribution at low mass is altered, which is demonstrated by the lami-1p-etairr model, which features a value of $\sigma$ consistent with that of an irradiated disk. Figure 7 illustrates that in this model low-mass bodies accrete a significant amount of pebbles. As $\sigma$ increases, this trend becomes more pronounced and results in poorer fits because the OSSOS++ part of the distribution can no longer be matched. This finding is unsurprising as the shape of the IMF was adopted from the cold classicals, which was already shown to fit the shape of hot classicals across the OSSOS++ range (Kavelaars et al., 2021; Petit et al., 2023). However, it also shows that accretion processes cannot maintain the shape of the low-mass part, at least not in conjunction also matching the high-mass tail. It lends further support to the primordial nature of the low-mass part of the TNO size distribution.
@@
An additional concern for these small $\tau_{s}$ runs is that they take a long time ($t/t_{0}$ is high). We can estimate the characteristic accretion time for pebble accretion from the pebble accretion rate (Equation (10)): $t\sim q/\dot{q}\sim t_{0}/\tau_{s}$, for $f_{\mathrm{set}}\sim 1$ (i.e., $m>m_{\ast}$). This estimate agrees with Table 3 where we compiled $t/t_{0}$ for selected model runs. We recall that the simulations are conducted in dimensionless parameter space, where we can hide our ignorance of, the pebble density, for example, which sets the value of $t_{0}$ (Equation 5). However, the pebble accretion mechanism requires gas-rich conditions and it must terminate before the lifetime of the gas-rich disk, typically estimated at ${\sim}5\,\mathrm{Myr}$ (Weiss et al., 2021). Therefore, a high $t/t_{0}$ renders these models implausible. We return to the timescale constraints in Section 4.
@@
Table 2 shows that a turbulent velocity field marginally improves the log-likelihood of the best model by an amount of +0.5. This specific run is characterized by a very low $\tau_{s}$ value (turb-1p-best entry in Table 3), which is disfavored in light of the above discussion. A higher $\tau_{s}$ run, following the $\tau_{s}$$\sigma$ correlation (turb-1p-hitau) fits the data equally well. Figure 6a presents the cumulative distribution of several individual runs. All models overshoot the infliction point at $m=10^{-4}\,m_{\oplus}$ and also tend to undershoot the final bin where the observed CMF terminates (with Triton) at $m=3\times 10^{-3}\,m_{\oplus}$. These apparent mismatches are a consequence of the low number of TNOs at these masses, which can be understood from inspecting the differential distribution (on which the likelihood calculations are based). Figure 6b presents these fits for the MPC part of the distribution. Here, the error bars denote Poisson counting uncertainties. The bin to the right of $m=10^{-4}\,m_{\oplus}$ with 0 bodies in the MPC is typically fitted with ${\approx}2$ bodies in the simulation, but this still has a Poisson likelihood of 13.5%. Similarly, the simulation runs are quite insensitive to the presence of the two most massive bodies (Pluto and Triton); statistically, the low number of high-$m$ bodies do not carry much weight. As a result the runs that undershoot the cumulative curve in Figure 6a (e.g., turb-1p-best) statistically cannot be distinguished from those runs that include a larger $\tau_{s}$ (e.g., turb-1p-hitau) and those runs that contain more pebbles (e.g., turp-1p-himass).
@@ 3.2 A second particle component
The addition of a secondary component does not significantly improve the results. From Table 2, the Bayes factor (the difference in $\ln\mathcal{P}$) is $0.54$ between the laminar models and $0.18$ between the turbulence models. These values indicate that the added parameters have insignificant effects on the fits (Jeffreys, 1998). This insensitivity is reflected in Table 2 by the proximity of the size of the secondary component ($\tau_{2}$) to that of the main component ($\tau_{s}$) in the lami-2p and turb-2p models. Figure 6 presents this result graphically. The lami-2p-best run—the run where $\mathcal{P}$ is the highest—hardly stands out from the other runs. Curiously, these models tend to be fitted by low $\tau_{s}$ (and $\tau_{2}$) and correspondingly high $t_{0}$, indicating that these runs improve the fit for the OSSOS++ part of the distribution, rather than the high-mass MPC part. In other words, the low number of high-mass bodies in P48 prevents any meaningful fit for the high-mass end of the distribution, which makes the model insensitive to a second particle component.
While we do not find evidence for a binary distribution of pebbles, two points should be noted. First, the size distribution could still be a power law, which would ameliorate the $f_{\mathrm{set}}$ transition in a similar way to that of the turbulence velocity field (Lyra et al., 2023). Second, the lack of evidence for a second component does not preclude its presence. A bigger sample of high-mass bodies is necessary to establish or discount its existence.